### v2

# RETTA: Retrieval-Enhanced Test-Time Adaptation for Zero-Shot Video Captioning

Abstract
Despite the significant progress of fully-supervised video captioning, zero-shot methods remain much less explored. In this paper, we propose a novel zero-shot video captioning framework named Retrieval-Enhanced Test-Time Adaptation (RETTA), which takes advantage of existing pre-trained large-scale vision and language models to directly generate captions with test-time adaptation. Specifically, we bridge video and text using four key models: a general video-text retrieval model XCLIP, a general image-text matching model CLIP, a text alignment model AnglE, and a text generation model GPT-2, due to their source-code availability. The main challenge is how to enable the text generation model to be sufficiently aware of the content in a given video so as to generate corresponding captions.To address this problem, we propose using learnable tokens as a communication medium among these four frozen models GPT-2, XCLIP, CLIP, and AnglE. Different from the conventional way that trains these tokens with training data, we propose to learn these tokens with soft targets of the inference data under several carefully crafted loss functions, which enable the tokens to absorb video information catered for GPT-2. This procedure can be efficiently done in just a few iterations (we use 16 iterations in the experiments) and does not require ground truth data. Extensive experimental results on three widely used datasets, MSR-VTT, MSVD, and VATEX, show absolute 5.1$\sim$32.4improvements in terms of the main metric CIDEr compared to several state-of-the-art zero-shot video captioning methods.
keywords:
[…]
1 Introduction
Video captioning, aiming to automatically describe the video content in one sentence, is one of the traditional vision-language tasks. It plays a crucial role in numerous applications, such as video title generation [1], blind assistance [2], and video search [3]. Although current video captioning research makes remarkable progress on full-supervision algorithms, exploration in the area of zero-shot learning still remains limited.
Recently, with the development of pre-trained foundation models in vision and language areas, their combination has shown outstanding zero-shot learning capabilities. For example, for the text-based question-answering task, pre-trained retrieval models have been integrated with Pre-trained Large language Models (PLMs) to improve the quality of generated answers. An example is shown in Fig. 1 (a), where the retriever takes the question “The 2024 Olympics will be held” as a query to retrieve relevant knowledge “The 2024 Summer Olympics, known as Paris 2024, will be hosted in Paris, France, from 26 July to 11 August.”. By concatenating it with the question as a prompt for the model, we get the output “Paris, France, from 26 July to 11 August 2024”.
Figure 1: Traditional retrieval-augmented-generation paradigm performs well on pure natural language tasks, such as text-based question answering. However, its performance sucks when directly apply to multi-modality tasks, such as video captioning.
However, a simple application of this paradigm does not work well for multi-modal visual description tasks (e.g., video captioning). For example, as shown in Fig. 1 (b), the video query retrieves the related sentence “Cutting a piece of bread”; then together with the hard prompt “Video showing” serving as the input of PLM, the PLM outputs “how to cut”, which does not meet the expectation of video captioning. The main reasons lie in two aspects: (1) Weak semantic connection: The semantic connection between the prompt “video showing” and the retrieved text is not as strong as that between questions and answers in QA tasks. This hampers PLMs to generate expected outputs. (2) Improper prompts: Prompt engineering itself is a challenging task, letting alone to additionally take diverse retrieved text into consideration.
To address the above-mentioned problems, we propose to leverage learnable tokens as a connection medium among different foundation models. These tokens should convey video information to PLMs and indicate to PLMs what type of output is expected. Moreover, we expect the combination of these pre-trained models is able to quickly adapt to the task, preferably without the need of training. To meet the quick adaptation requirement, we propose to update these tokens during inference but with soft targets, which gets rid of the conventional training of soft prompts. Our experiments show that the whole framework can well adapt to the video captioning task after just a few of iterations (we use 16 iterations). To meet the former content awareness and output type requirements, we design several loss functions aligning the generated text and soft targets. Specifically, the soft targets are calculated by retrieved text and candidate sentences to guide the choice of candidate words at each time step. This is because soft targets can help the model consider multiple potential tokens in text generation, boosting the diversity of the generated results and improving generalization. An example is shown in Fig. 1(c), where our method leads to better video captions.
In summary, the main contributions of this paper are as follows:
  • To the best of our knowledge, we propose the first fast adaptation framework that combines a pre-trained video-text retrieval model, an image-text matching model, a text alignment model, and a language generation model for video captioning.
  • We design two loss functions that effectively convey retrieved information to PLMs to generate captions closely related to video content.
  • Extensive experimental results on three widely-used datasets show that our method outperforms several state-of-the-art zero-shot video captioning approaches with absolute 5.1$\sim$32.4improvement in terms of the main metric CIDEr.
2 Related work
Video Captioning Recently, a series of full-supervision video captioning research [4, 5, 6, 7, 8, 9, 10, 11, 12] have made significant progress. However, fewer efforts have been made on the video captioning task under the zero-shot setting [13, 14, 15, 16]. ZeroCap [13] uses a visual-language model to guide a language model to describe the given image via the designed CLIP-loss. EPT [14] further extends it to video captioning using the proposed CLIP-loss to optimize the pseudo-tokens. To improve inference speed, MAGIC [15] optimizes the decoding strategy of language models and employs CLIP scores to adjust PLM logits for better alignment with image correspondence. DeCap [17] and MultiCapCLIP [18] also leverage the visual-language alignment ability of CLIP for zero-shot video captioning. They first utilize text-only data to train a decoder from scratch and then input video to generate captions at the inference stage. Although the above-mentioned methods improve zero-shot caption quality, they still face the challenge of understanding video content. To tackle this problem, we propose a retrieval-augmented video captioning framework that retrieves video-related knowledge to deepen the understanding of the video.
Video-Text Retrieval Video-text retrieval (VTR) aims to find the most relevant text based on a video query. To retrieve the correct video candidates given a text query, recent works [19, 20, 21, 22, 23, 24] focus on the cross-modal alignment between the related videos and text samples. Most of them can be classified into coarse-grained alignment or fine-grained alignment. The coarse-grained alignment methods mainly leverage the frame-level or video-level visual features to perform alignment between the video and query features. Clip4clip [19] transfers the image-text knowledge of the CLIP model to the VTR task in a coarse manner. X-Pool [20] designs a cross-modal attention model that performs coarse-grained alignment between a text and the frames of a video. TS2-Net [21] introduces the token shift and selection transformer to better model video representation in both temporal and spatial dimensions. However, the coarse-grained alignment is not suitable for capturing detailed correspondence between texts and videos. To overcome this challenge, TokenFlow [22] presents a fine-grained alignment for token-wise similarity calculation. X-CLIP [23] adopts multi-grained contrastive learning to boost video-text retrieval performance, covering fine-grained, coarse-grained, and cross-grained contrast. Another X-CLIP [24] extends CLIP for general video-language understanding using the Kinetics-600 [25] dataset. In this work, we adopt [24] as our retriever as its training data has no overlap with video captioning datasets.
Retrieval-Augmented Generation Retrieval-Augmented Generation (RAG) augments the generation ability of language models with retrieved knowledge. Generally, the knowledge retrieved is used as a prefix input to the pre-trained language model (PLM), prompting PLM to generate more accurate answers. REALM [26] employs a latent knowledge retriever to boost language model pre-training, which allows capturing more real-world knowledge in an interpretable manner. ORQA [27] pre-trains the retriever with an unsupervised inverse cloze task and learns to retrieve evidence from an open corpus. DPR [28] implements a dense embedding retriever to better select candidate contexts using only pairs of questions and answers, without additional pretraining. RCG [29] plugs a cross-modal retriever to provide useful hints and employs a dynamic copy-mechanism generator for better generation. Recently, with the development of large language models (LLMs), a series of RAG-related LLM approaches are proposed. RECOMP [30] uses a summarization model to compress retrieved documents into a textual summary before prompting the LLM to generate the output. Chen et al. establish the first benchmark to evaluate robustness-related capabilities for retrieval augmented generation of LLMs. Asai et al. further introduce a new framework SELF-RAG, which trains an arbitrary LM to generate text with reflection tokens by integrating them as the next token prediction from the expanded model vocabulary.
Although RAG has been applied to fully-supervised video captioning, there is no related work on zero-shot video captioning tasks, to the best of our knowledge. It faces the challenge of how retrieved sentences can guide the LLMs in generating appropriate video descriptions. To solve this problem, we design two loss functions to optimize learnable tokens via aligning LLMs’ output and retrieval results.
Figure 2: The workflow of the proposed method, which consists of frozen pre-trained foundation vision and language models as well as trainable tokens. Unlike conventional soft prompt training, we update these tokens directly during inference with soft target. In this way, the frozen foundation models can quickly adapt to the video captioning task for zero-shot application. $\mathcal{L}_{language}$ is computed between token probability distributions generated with (i.e. writing) and without (i.e. reading) the trainable soft prompt. $\mathcal{L}_{vision}$ calculates a matching score between the current generated text and visual information leveraging CLIP based on keyframes. $\mathcal{L}_{retrieval}^{S}$ measures a matching score between the current generated text and retrieved sentences using AnglE, and $\mathcal{L}_{retrieval}^{W}$ focuses on the high-frequency words of them. The above optimizations take place at each step of the autoregressive process.
3 Method
As shown in Fig. 2, our framework contains four pre-trained frozen models: XCLIP [24] and CLIP [31] provide video-level and frame-level video understanding; GPT-2 [32] provide the text generation capability; AnglE [33] is used to calculate text-text similarity. According to their functions for the video captioning task, these pre-trained models can be grouped into a video-text retrieval module and an augmented caption generation module. Specifically, XCLIP serves as the video-text retrieval module, providing a basic understanding of the video. It retrieves video content-related sentences from training text corpus, and we highlight its high-frequency words, as detailed in Section 3.1. The remaining pre-trained models compose the augment caption generation module, which utilizes several test-time learnable tokens to connect all these pre-trained models, absorb video information, and serve as proper prompts for text generation as detailed in Section 3.2. In Section 3.3, we provide more details of the inference process.
3.1 Video-Text Retrieval
The video-text retrieval module adopts a multi-modal retriever XCLIP to retrieve video-related descriptions from a text corpus. XCLIP is pre-trained on the Kinetics-600 [25] dataset and has the zero-shot ability for video classification and video-text retrieval tasks. Traditional retrievers consistently fetch the top-$k$ ranked sentences. However, this retrieval strategy faces two challenges for video captioning. First, choosing the hyper-parameter $k$ is challenging since the optimal $k$ varies across different videos. Second, retrieving a fixed number of $k$ sentences may introduce noise, as not all of these sentences are closely related to the video.
To mitigate the aforementioned issues, we propose to additionally sample top-$L$ high-frequency words (limited to nouns and verbs) from the retrieved sentences, which reduces diversity of sentences. We detail how retrieved sentences and sampled high-frequency words are utilized to guide caption generation in the next section.
3.2 Augmented Caption Generation
With the help of the video-text retrieval module, the augmented caption generation module uses the retrieved sentences and high-frequency words to improve the quality of the generated captions. This module contains three main components: the large language model GPT-2, the image-text matching model CLIP, and the text alignment model AnglE. The out-of-box GPT-2 is pre-trained on the large dataset WebText collected from a wide range of text data online. It is composed of several Transformer layers, which can cache the previous key and value tokens to infer the next token autoregressively. Therefore, GPT-2 is capable of generating text based on preceding content. The CLIP model is pre-trained on 400 million image-text pairs, which plays the vision-language alignment role to map the image and sentence into a unified semantic space. AnglE is a state-of-the-art model for calculating semantic textual similarity, which is able to be aware of differences among texts.
As shown in Fig. 2, all model parameters are frozen, and only the learnable soft prompts are optimized during the inference process with soft targets to guide GPT-2 to generate text in line with the content of the given video. Take the $i-{th}$ step as an example, the text generation can be described as:
$w_{i+1}=\operatorname{PLM}(w_{soft},w_{hard},s_{i})$ (1)
where $w_{i+1}$ denotes the next to be generated word and $s_{i}$ denotes all the existing words generated in previous steps. $w_{soft}$ is the set of learnable tokens that can be updated by the guidance during the generation process. $w_{hard}$ is the hard prompt, which is randomly sampled from a predefined prompt set, such as “Video showing”, “Video of”.
To utilize the retrieved sentences, we transfer them into the above auto-regression process with our proposed retrieval-based loss $\mathcal{L}_{retrieval}$. It consists of two granularities: the sentence-granularity $\mathcal{L}_{retrieval}^{S}$ and the word-granularity $\mathcal{L}_{retrieval}^{W}$ Specifically, the $\mathcal{L}_{retrieval}^{S}$ is formulated as
${}\mathcal{L}_{retrieval}^{S}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{p}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (2)
where $\mathrm{q}\left(\mathrm{w}_{k}\right)$ is the probability of the $k$-th word in the vocabulary, and $\mathrm{p}(\cdot)$ denotes the probability of being soft target. $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is computed as
$\mathrm{p}\left(\mathrm{w}_{k}\right)=\frac{1}{K}\sum_{R\in\mathcal{R}}\operatorname{AnglE}\left(R,s_{k}\right)$ (3)
where $K$ denotes the number of retrieved sentences, $s_{k}$ represents the combination of the generated words in previous time steps with the $k$-th word, $\mathrm{p}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the retrieved sentence in $\mathcal{R}$. This loss adjusts learnable tokens so as to prompt GPT-2 to generate texts better reflect the content in the retrieved description of the given video. For efficiency, we only calculate the scores for the top 100 candidate words based on the original probability distribution obtained by GPT-2.
On the other hand, considering that the retrieved sentences might contain noise, we propose a word-granularity loss, which focuses on the high-frequency information in the retrieved sentences. The word-granularity retrieval loss is formulated as
${}\mathcal{L}_{retrieval}^{W}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (4)
$\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)=\frac{1}{L}\sum_{W\in\mathcal{W}}\operatorname{AnglE}\left(W,s_{k}\right)$ (5)
where $\mathcal{W}$ denotes all the high-frequency words in the retrieved sentences. $\mathrm{\hat{p}}\left(\mathrm{w}_{k}\right)$ is the average of $\operatorname{AnglE}$’s alignment scores between $s_{k}$ and all the sampled high-frequency words in $\mathcal{W}$.
Inspired by [13, 14, 16], we also employ two reliable losses to optimize the augment caption generation module. The first loss aims to reduce the gap between the generated caption and the corresponding video:
$\mathcal{L}_{vision}=-\sum_{\mathrm{k}=1}^{\mathrm{N}}(\frac{1}{T}\sum_{F\in\mathcal{F}}\operatorname{CLIP}\left(F,s_{k}\right))\log\left(\mathrm{q}\left(\mathrm{w}_{k}\right)\right)$ (6)
where $T$ is the total number of sampled video frames. CLIP calculates the average matching score between each sampled frame $F$ in $\mathcal{F}$ and $s_{k}$.
The second loss maintains the natural language fluency of the generated caption while aligning it with the video content:
$\mathcal{L}_{\text{language}}=\text{CE}\left(\mathrm{q}(w_{i+1}),\mathrm{q}(\hat{w}_{i+1})\right)$ (7)
$\hat{w}_{i+1}=\operatorname{PLM}(w_{hard},s_{i})$ (8)
where CE is the cross-entropy loss function, $\hat{w}_{i+1}$ is the distribution of generated token without soft prompt.
3.3 Inference
In our zero-shot setting, we directly perform test-time learning, where only text corpus of training split are used for retrieving caption candidates. The retrieved captions are used to generated soft targets as mentioned in Sec. 3.2. The learnable tokens are optimized by the total loss:
$\mathcal{L}_{total}=\lambda_{l}\mathcal{L}_{language}+\lambda_{s}\mathcal{L}_{retrieval}^{S}+\lambda_{w}\mathcal{L}_{retrieval}^{W}+\lambda_{v}\mathcal{L}_{vision}$ (9)
where $\lambda_{l}$, $\lambda_{s}$, $\lambda_{w}$ and $\lambda_{v}$ are the corresponding loss weights.
Optimization occurs throughout the sentence generation, enhancing the correspondence between the generated sentence and video through the iterative application. More specifically, the optimization is conducted at each step in the auto-regression process. When generating the stop token or reaching the maximum sentence generation length, we stop the process and initiate the generation of a new sentence while retaining the optimized soft prompt.
Figure 3: Our sentence-cleaning post-process strategy extract key elements, and use them to reconstruct the clean caption with the pre-trained CommonGen.
After all $M$ iterations are complete, we obtain $M$ generated captions. We calculate the similarity between these $M$ generated captions and the video as follows:
$s^{*}=\arg\max_{s_{m}\in\{s_{1},s_{2},\ldots,s_{M}\}}\left(\frac{1}{T}\sum_{F\in\mathcal{F}}\operatorname{CLIP}(F,s_{m})\right).$ (10)
The caption $s^{*}$ with the highest similarity is selected as the reliable caption. However, due to the limitations of zero-shot learning, the selected caption still contains some noise and lacks fluency. We design a simple but effective sentence-cleaning strategy, which consists of two steps: (1) extracting essential elements like verbs and nouns from the selected sentence, and (2) employing a pre-trained keyword-sentence model CommonGen [34] to reconstruct these elements into a well-organized sentence. CommonGen is a state-of-the-art model for constrained text generation, which takes keywords (such as “knife”, “chopping”, “onion” and “kitchen”) as input and generates a more fluent sentence “a man is chopping onions with a knife in a kitchen”, as shown in Fig. 3. In the end, the cleaned sentence is used as the final result.
4 Experiments
4.1 Datasets and Metrics
MSR-VTT [35] is a large-scale video captioning dataset collected from a commercial video website. It contains 10,000 video clips covering 20 categories. Each clip has 20 English annotations. In the standard split, it includes 6,513 training videos, 497 validation videos, and 2,990 test videos.
MSVD [36] is a widely-use video captioning dataset, which is collected from Youtube. It consists of 1,970 video clips. Each clip has roughly 40 ground-truth captions. In the standard split, it includes 1,200 training videos, 100 validation videos, and 670 test videos.
VATEX [37] is a large-scale dataset that contains 34,991 videos with 10 English annotations. The standard split includes 25,910 training videos, 3,000 validation videos, and 6,000 test videos.
Metrics. In our experiments, we adopt four commonly used metrics to evaluate caption quality, including BLEU@4 (B@4) [38], METEOR (M) [39], ROUGE-L (R) [40], and CIDEr (C) [41]. Among these metrics, CIDEr is the main metric.
4.2 Implementation Details
We use the pre-trained XCLIP [^1] https://huggingface.co/microsoft/xclip-base-patch16-kinetics-600-16-frames as our video retriever. We uniformly sample 16 frames per video for video retrieval. The visual feature extraction and text-image similarity calculation are performed using the pre-trained CLIP ViT-L/14222https://github.com/openai/CLIP model. We first sample three frames every second for the given video. To avoid using duplicate frames, we further use the CLIP-based strategy to subsample the keyframes. We start with the first frame as an anchor. The next frame is included in the set $\mathcal{F}$ if its dot product with the anchor, calculated by the CLIP image encoder, exceeds a specific threshold $\lambda_{CLIP}$. The newly sampled frame becomes the new anchor and the sampling process repeats in the same manner. The $\lambda_{CLIP}$ is set to 0.9 in all experiments. We use the pre-trained GPT-2 medium [^3] https://huggingface.co/openai-community/gpt2-medium as our caption generator, and the pre-trained AngLE [^4] https://github.com/SeanLee97/AnglE to calculate the similarity between the generated caption and retrieval sentences. Last, we leverage the pre-trained CommonGen [^5] https://huggingface.co/mrm8488/t5-base-finetuned-common_gen to clean the generated sentences into fluent ones. The aforementioned pre-trained models are publicly and readily available.
Empirically, $\lambda_{l}$, $\lambda_{v}$, $\lambda_{s}$ and $\lambda_{w}$ are set to 1.6, 1.0, 0.8, and 0.3 respectively. In our method, we set the number of learnable tokens to 5. We set the number of retrieval sentences to 15 and the frequency size to 5. We adopt the AdamW [42] optimizer with a learning rate of 1e-4 and a weight decay weight of 0.3 to optimize our model. Throughout the experiments, we employ 16 generation iterations per video. To prevent sentences from being overly long and repetitive, we limit the number of tokens generated per sentence to 15. We implement our framework with PyTorch and all experiments are conducted on 1 $\times$ RTX-A6000 GPU.
Table 1: Results on MSVD across various models and supervision levels.
Methods Setting Metrics
B@4 M R C
Results on MSVD test set
VNS-GRU [4] Supervised 66.5 42.1 79.7 121.5
SemSynAN [5] 64.4 41.9 79.5 111.5
HMN [43] 59.2 37.7 75.1 104.0
ZeroCap [13] Zero-shot 2.9 16.3 35.4 9.6
MAGIC [15] 6.6 16.1 40.1 14.0
EPT[14] 3.0 17.8 31.4 17.4
Video-LLaMA[44] 4.0 15.1 21.6 1.3
RETTA(Ours) 23.3 28.5 56.4 49.8
Table 2: Results on MSRVTT across various models and supervision levels.
Methods Setting Metrics
B@4 M R C
Results on MSR-VTT test set
VNS-GRU [4] Supervised 45.3 29.9 63.4 53.0
SemSynAN [5] 46.4 30.4 64.7 51.9
HMN [43] 43.5 29.0 62.7 51.5
SGN [45] 40.8 28.3 60.8 49.5
SAAT [46] 40.5 28.2 60.9 49.1
ZeroCap [13] Zero-shot 2.3 12.9 30.4 5.8
MAGIC [15] 5.5 13.3 35.4 7.4
EPT[14] 3.0 14.6 27.7 11.3
Video-LLaMA[44] 4.9 16.8 25.3 2.3
ZS-CapCLIP[31] 4.0 15.0 31.0 5.0
MultiCapCLIP[18] 13.3 19.5 43.3 15.5
DeCap-BookCorpus [17] 6.0 12.7 - 12.3
DeCap-CC3M [17] 6.2 14.9 - 15.0
DeCap-COCO [17] 14.7 20.4 - 18.6
RETTA(Ours) 14.0 19.3 42.2 24.3
Table 3: Results on VATEX across various models and supervision levels.
Methods Setting Metrics
B@4 M R C
Results on VATEX public test set
VaTeX [37] Supervised 28.4 21.7 - 45.1
HMN [43] 33.2 22.7 49.3 51.6
DeCap-BookCorpus [17] Zero-shot 4.1 9.9 - 11.8
DeCap-CC3M [17] 7.3 12.6 - 18.4
DeCap-COCO [17] 13.1 15.3 - 18.7
Video-LLaMA[44] 4.3 16.3 21.8 3.8
RETTA(Ours) 11.4 16.3 32.6 23.8
4.3 Performance Comparison
We compare our method with several state-of-the-art models under different settings on three public benchmark datasets. The results are shown in Table 1, 2 and 3. We note that the performance of supervised algorithms is significantly higher than zero-shot methods, but they require expensive annotation costs. Compared to the other zero-shot methods, our method achieves the best performance in terms of the main metric CIDEr on all the three datasets. As shown in Table 1, our method achieves a CIDEr score of 49.8 on the MSVD benchmark, leading by about an absolute 32.4 over the previous best result. On the more challenging benchmark MSR-VTT, our method also boosts the CIDEr score by about 30.6%. On the VATEX dataset, our method achieves a 23.8 CIDEr Score, gaining 27.3% improvement in terms of the main metric CIDEr on VATEX.
4.4 Ablation Study
In this section, we perform a set of ablation studies on VATEX as it is the largest and most challenging.
Table 4: Performance with different our proposed component. Experiments are conducted on the VATEX dataset.
# $\mathcal{L}_{retrieval}^{S}$ $\mathcal{L}_{retrieval}^{W}$ $\mathcal{S}_{clean}$ B@4 M R C
1 1.4 9.8 19.4 8.9
2 2.5 13.3 22.6 20.8
3 2.6 12.2 22.1 14.9
4 2.4 13.5 22.9 21.4
5 11.4 16.3 32.6 23.8
Effect of our proposed component. To fully analyze the effect of each proposed component, we conduct experiments on the challenge VATEX. As shown in Table 4, we observe that each component assists the baseline to get better performance on all metrics. For example, Row 2 shows that using only sentence-granularity retrieval loss leads to a CIDEr score increase of $133.7\%$ compared to the baseline in Row 1. In contrast, employing word-granularity retrieval loss results in a CIDEr enhancement of $67.4\%$ and a BELU@4 score improvement of $85.7\%$ (Row 3 v.s. Row 1). The second-best performance is achieved when both types of losses are used together (Row 4 v.s. Row 1). In the comparison, the results with $\mathcal{S}_{clean}$ post-processing reach a further improvement of $11.2\%$ in the term of CIDEr (Row5 v.s.  Row4).
Effect of different retrieval corpora. In Table 5, we compare the performance of our method under different retrieval corpora. On the MSR-VTT benchmark, the CIDEr score increases around $9.9\%$ by leveraging the TestSet as the retrieval corpus instead of the TrainSet (Row 1 v.s. Row 2). On MSVD and VATEX datasets, employing TestSet improves around $21.9\%$ (Row 3 v.s. Row 4) and $8.4\%$ (Row 5 v.s. Row 6) compared to TrainSet. This demonstrates that different knowledge corpora can significantly impact performance, especially on relatively small datasets (e.g., MSVD). In this paper, unless otherwise specified, we use TrainSet as our default text corpus.
Effect of video retriever. In order to study the impact of video retrievers on our method, we use different configurations of XCLIP as our retriever. The results are present in Table 6. We observe that when XCLIP utilizes a ViT [47] with a size of B/16 as the backbone, it achieves an approximate improvement of 2.2% in the CIDEr score compared to using B/32 (Row 2 v.s. Row 1). The Kinetics-600 [25] dataset is the extended version of Kinetics-400 [48], containing more training data. Therefore, XCLIP pre-trained on Kinetics-600 brings a further improvement of 1.7% compared to XCLIP pre-trained on Kinetics-400 (Row 3 v.s. Row 2). Row 4 achieves the best performance on all metrics. It is trained with 16 frames per video and outperforms Row 3 using 8 frames. Therefore, in the other experiments conducted in this paper, we use the XCLIP in Row 4 as our default retriever.
Effect of the number of retrieved sentences. We study the effect of using different numbers of retrieved sentences. Figure 4 shows the difference in video captioning performance as $K$ varies from 5 to 20. We find that retrieving 15 sentences offers better CIDEr scores and BLEU@4 scores compared to other numbers. It can be concluded that too many or too few retrieval sentences can degrade the performance. This is because too many retrieved sentences may introduce too much noise while too few retrieved sentences cannot provide sufficient information.
Table 5: Effect of retrieval corpora. The experiments are conducted on test split of MSVD, MSR-VTT and VATEX datasets.
Dataset # Retrieval Corpus B@4 M R C
MSR-VTT 1 TrainSet 14.0 19.3 42.2 24.3
2 TestSet 15.3 20.2 43.1 26.7
MSVD 3 TrainSet 23.3 28.5 56.4 49.8
4 TestSet 29.4 31.8 60.0 60.7
VATEX 5 TrainSet 11.4 16.3 32.6 23.8
6 TestSet 12.3 17.1 33.2 25.8
Table 6: Performance with different configuration XCLIP. The experiments are conducted on the VATEX dataset.
# Backone $N$ Pre-train B@4 M R C
1 VIT-B/32 8 Kinetics-400 10.5 15.6 31.9 22.3
2 VIT-B/16 8 Kinetics-400 10.9 15.8 32.2 22.8
3 VIT-B/16 8 Kinetics-600 11.2 16.1 32.3 23.2
4 VIT-B/16 16 Kinetics-600 11.4 16.3 32.6 23.8
Effect of the number of learnable tokens. We study the effect of the number of learnable tokens, and Figure 5 shows that 5 is the better choice than other numbers. In Figure 5(a), the CIDEr score initially decreases as the number increases from 1 to 3, then rises to its highest point from 3 to 5, followed by another decrease from 5 to 7. In Figure 5(b), the Bleu@4 score first increases as the number goes from 1 to 5, then decreases from 5 to 7. Therefore, we use 5 as our default number of soft tokens.
(a) CIDEr under different numbers of retrieved sentences
(b) BLEU@4 under different numbers of retrieved sentences
Figure 4: Performance (CIDEr and BLEU@4 score) comparison under different numbers of retrieved sentences on the VATEX dataset.
(a) CIDEr under different numbers of soft tokens
(b) BLEU@4 under different numbers of soft tokens
Figure 5: Performance (CIDEr and BLEU@4 score) comparison under different numbers of soft tokens on the VATEX dataset.
Effect of the frequency size of the word-granularity loss. In Fig. 6, we conduct experiments on VATEX to explore the impact of the frequency size $L$ of word-granularity loss. From the figure, we observe that the CIDEr score first rises before reaching the moderate number (i.e., $L=5$), and then starts to decrease slightly. The main reason is that when $L$ is large, some noisy words are introduced. On the contrary, if $L$ is small, some essential words may be neglected. In Fig. 6(b), a similar trend is observed in the BLEU metric as well. Therefore, we select an appropriate number (i.e., $L=5$) as the frequency size.
(a) CIDEr under different frequency size
(b) BLEU@4 under different frequency size
Figure 6: Performance (CIDEr and BLEU@4 score) comparison under different frequency sizes on the VATEX dataset.
Effect of the number of candidate words As mentioned in Sec. 3.2, to enhance efficiency, we compute the loss using the top 100 candidate words based on the original probability distribution obtained by the language model. In Figure 7(a) and Figure 7(b), we conduct experiments on VATEX to study the impact of the number of candidate words. From Figure 7(a), we find that using the top 250 candidate tokens achieves the highest CIDEr score compared to other numbers, with the top 100 candidate tokens achieving the second-best CIDEr score. In contrast, in Figure 7(b), we observe that the average time consumed per video increases roughly linearly with the number of candidate tokens. Compared to using 100 tokens, using 250 tokens takes about twice as long. To balance performance and time consumption, we set the number of candidate words to 100.
(a) CIDEr under different numbers of candidate words
(b) Average runtime under different numbers of candidate words
Figure 7: CIDEr and Average runtime per video under different numbers of candidate words on the VATEX dataset.
Effect of hard prompt set In our method, we randomly sample a phrase from a predefined set as the hard prompt. The set we designed includes four phrases: “Video showing”, “Video of”, “Video shows” and “Video describes”. In Table 7, we compare the performance of our method under different sizes of hard prompt sets. Row 1 displays the worst CIDEr score in the absence of any hard prompts. Using “Video showing” as a hard prompt consistently results in a CIDEr improvement of about 1.7% (Row 2 v.s. Row 1). The inclusion of “Video showing” and “Video of” in the hard prompt set leads to a CIDEr increase of around 0.8% (Row 3 v.s. Row 2). With the addition of “Video shows”, the CIDEr score in Row 4 leads by an absolute 0.3 over Row 3. Therefore, in our other experiments, the phrases covered in Row 4 are set as our default hard prompt set.
Table 7: Performance with different sizes of hard prompt sets. Experiments are conducted on VATEX.
# Hard Prompt Metrics
Video Video Video B@4 M R C
showing of shows
1 10.7 16.0 32.0 22.9
2 10.9 16.1 32.3 23.3
3 11.2 16.4 32.8 23.5
4 11.4 16.3 32.6 23.8
Zero-shot image captioning. Our method can be adapted for zero-shot image captioning by swapping XCLIP with CLIP. In Table 8, we compare the performance among several zero-shot image captions methods on the MS-COCO [49] dataset. We achieve the second-best results on all metrics, only falling behind MAGIC [15]. This is because MAGIC fine-tuned the language model on the text corpus of MS-COCO captions. Compared to ZeroCap [13] and EPT [14], our CIDEr score improves by an absolute 31.9 and 27.8, respectively.
Table 8: Image captioning results on MS-COCO.
Methods Setting Metrics
B@4 M R C
Results on MS-COCO test set
ZeroCap [13] Zero-shot 2.9 12.0 - 13.1
MAGIC [15] 12.9 17.4 - 49.3
EPT[14] 2.2 12.7 - 17.2
RETTA(Ours) 11.8 19.7 35.9 45.0
4.5 Qualitative Results
Several video captioning qualitative results are shown in Fig. 8. We can see that the caption produced by our method is more specific than other methods. For example, the video in Fig. 8 (a) shows two men playing ping pong. Our method accurately describes “The two boys playing ping pong in a table tennis court”. In Fig. 8 (b), although the fully supervised method HMN describes “A man is riding a motorcycle”, our method more accurately identifies the “ocean wave”. In Fig. 8 (c), our method effectively describes “The tutorial instructions for downloading songs”, whereas the compared zero-shot method EPT merely mentions “download” but does not capture “songs”. In Fig. 8 (d), only our method correctly describes the video content “Cooking shrimp in oil”. In contrast, the generated caption “A person is preparing a recipe” by HMN is too general. These results further show the zero-shot ability of our proposed method.
Fig. 9 also shows several zero-shot image captioning qualitative results. In Fig. 9 (a), our method identifies “laptop computer” and correctly describes the image “A baby typing on laptop computer”. In Fig. 9 (b), while three methods all recognize “pizza”, our method additionally captures “plate”. In Fig. 9 (d), our method vividly generates “Elephant slide in the park”.
Figure 8: Several video captioning qualitative examples using different methods.
Figure 9: Several image captioning qualitative examples using different methods.
5 Conclusion
In this paper, we make the first attempt to explore the combination of a video-text retrieval model, an image-text matching model, and a large language model under the zero-shot and fast adaptation setting for video captioning. Specifically, we use learnable tokens as an information medium connecting these frozen models. With the help of our carefully designed loss functions, these tokens are able to convey video information to the LLM in a way that serves as suitable soft prompts. The soft target-based inference-time token update ensures fast adaptation. Our proposed framework is conceptually simple and easy to implement. The extensive experiments on three benchmarks demonstrate its favorable zero-shot ability. We will make the code publicly available and hope our work can inspire future research in other areas.